OpenStudy (anonymous):

What is the area of a sector with a central angle of 210 degrees and a diameter of 4.6 m?

OpenStudy (anonymous):

 The first thing I'd do is find the area of the entire circle. Area of Circle: pi*(radius)^2 = pi*( (4.6)/2) )^2 = pi*(2.3)^2 = 16.619 M. This is our whole area. This is good information...Let's see what we can do with it. We're looking to find the area of the 210 degree sector, so here's what we'll do. Make a fraction. (210 degrees) / (360 degrees) = (7/12) of the entire circle. So we know the (7/12) of the total circle, which has an area of 16.619 M, is equal to the area of the 210 degree sector. Let's solve that out. (7/12)(16.619 M) = 9.694 M. That's the area of the sector. Hope I helped! ^_^ PS: It's great that you trust in the Lord. Few people really do! Rely on Him every day, and He'll bless you richly, beyond your wildest dreams.